OpenStudy (anonymous):

can someone please help me? Medal awarded xx please!! evaluate no calculators 81^-3/4 (9/4)^-5/2 (1000/343)^4/3

OpenStudy (anonymous):

instead of (1000/343) the other problem is (27/1000)^-2/3

OpenStudy (whpalmer4):

\[81^{-3/4} = \frac{1}{81^{3/4}}\]by \[a^{-n} = \frac{1}{a^n}\]then by rules for fractional exponents, \[(x^a)^b = x^{a*b}\]so\[\frac{1}{81^{3/4}} = \frac{1}{(81^{1/4})^3}= \frac{1}{(\sqrt[4]{81})^3}\] Can you find the 4th root of 81?

OpenStudy (whpalmer4):

Hint:\[\sqrt[4]{x} = \sqrt{\sqrt{x}}\]

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

okay, so that leaves us with \[\frac{1}{(\sqrt[4]{81})^3} = \frac{1}{3^3} =\]

OpenStudy (anonymous):

1/9

OpenStudy (whpalmer4):

look more closely

OpenStudy (anonymous):

oh 27 haha

OpenStudy (whpalmer4):

much better. the others should be very similar...

OpenStudy (anonymous):

so looking at (27/1000)^-2/3, would i have to multiply by 3/2?

OpenStudy (whpalmer4):

\[(\frac{27}{1000})^{-2/3} = (\frac{1000 } {27})^{2/3}\]

OpenStudy (anonymous):

would the answer be 100/9?

OpenStudy (whpalmer4):

It would indeed.

OpenStudy (anonymous):

thank you soooo much! That makes so much sense now. Thank you for your time xx

OpenStudy (whpalmer4):

very easy to compute those if you realize that \(1000 = 10^3\) and \(27 = 3^3\), then it's just \[(\frac{10^3}{3^3})^{2/3} = \frac{10^{3*(2/3)}}{3^{3*(2/3)}}=\frac{10^2}{3^2}\]

OpenStudy (anonymous):

yes so for the (9/4)^-5/2 would that be the same concept?

OpenStudy (whpalmer4):

you could do it that way, yes. make the numbers jump! :-)

OpenStudy (anonymous):

:D hahahaha so basically it would be 4/9^5/2 format then solve it from there

OpenStudy (whpalmer4):

yes. or if you're doing it that last way, you could just wait until the end to flip it over. 6 of one, half dozen of the other...

OpenStudy (anonymous):

it would equal 32/243

OpenStudy (whpalmer4):

time for me to call it a night. looks like you have this well under control!

OpenStudy (anonymous):

thanks again so much! :) you are the best :D